# The modified diagonal cycles of Hypergeometric curves

### v1

@@ 3.2. Properties of the cusps
Our goal in this subsection is to establish some results about the cusp points. These results will be needed in the proof of Theorem 1.1, which will be given in §4. In particular, Lemma 3.5 will play a key role.
We start by the analogue of the Manin-Drinfeld theorem for hypergeometric curves:
@@
Let $P$ and $Q$ be cusps on $X_{N,\lambda}$. Then the point $P]-[Q]\in\operatorname{Jac}(X_{N,\lambda})$ is a torsion point of order dividing $N^{2}$. If $\{P,Q\}$ is any of $\{b_{i},c_{2,j}\}$, $\{a_{i},c_{1,j}\}$, $\{a_{i},a_{j}\}$, $\{b_{i},b_{j}\}$, $\{c_{1,i},c_{1,j}\}$ or $\{c_{2,i},c_{2,j}\}$ for any $i,j$ then [P][Q][P]-[Q] is annihilated by $N$.
@@
Let $K(X_{N,\lambda})$ be the canonical divisor of $X_{N,\lambda}$. For any cusp $e\in X_{N,\lambda}$, we have
$(2(N-1)^{2}-2)[e]=K(X_{N,\lambda})$
@@
Thus by Proposition 3.3, $K(X_{N,\lambda})-(2(N-1)^{2}-2)[e]$ is torsion in the Jacobian, which proves the assertion. ∎
Our final result in this subsection is a result about the space of meromorphic functions with bounded poles at the cusps. The result may seem rather unmotivated by itself, but it will play a key role in the next section. For a positive integer $d$ and points $e_{0},\ldots,e_{n}\in X_{N,\lambda}$, let
@@
be the vector space (over the constants $\mathbb{C}$) of functions on $X_{N,\lambda}$ that are regular outside $\{e_{i}\mid i=0,\ldots,n\}$ and do not have a pole of order greater than $d$ at any $e_{i}$.
@@
Let $p\neq 2$ be a prime number and $l$ a positive integer such that $l\leq(p-1)/2$. Then for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and every cusp $e\in X_{p,\lambda}$, the point
$l\cdot(\varphi^{a,b}_{p})_{*}([P]+[Q]-2[e])\in\operatorname{Jac}(C_{p,\lambda}^{a,b})$
@@
We only prove the case $e=c_{1,0}$ since the other cases are proved similarly. It suffices to show that the point
$l\cdot(\varphi_{p}^{a,b})^{*}\circ(\varphi^{a,b}_{p})_{*}([P]+[Q]-2[e])=l\cdot\sum_{j=0}^{p-1}\left(g_{p}^{bj,-aj}\cdot([P]+[Q])-2c_{1,j}\right)$
of $\operatorname{Jac}(X_{p,\lambda})$ is nontrivial. Let $K$ be the function field of $X_{p,\lambda}$. Suppose that the above point of $\operatorname{Jac}(X_{p,\lambda})$ is trivial. Then there exists a function $f\in K$ such that
$\operatorname{div}(f)=l\cdot\sum_{j=0}^{p-1}\left(g_{p}^{bj,-aj}\cdot([P]+[Q])-2c_{1,j}\right).$
@@
On the other hand, $g$ is a polynomial in only $x$ and therefore $g$ does not have any poles outside $c_{1,i}$ ($i=0,\ldots,p-1$), contradicting (4.1). ∎
We will now deduce Theorem 1.1. As in the statement of the theorem, suppose $p$ is a prime number $\geq 3$. For brevity, we set $X=X_{p,\lambda}$. Let $e$ be a cusp of $X$. We take $\Gamma_{\alpha}\in\operatorname{CH}_{1}(X\times X)$ to be the graph of $\alpha$. Let $\operatorname{pr}_{123}\colon X^{3}\times X\to X^{3}$ (resp. $\operatorname{pr}_{4}\colon X^{3}\times X\to X$) be the natural projection to the first three factors of $X$ (resp. the fourth factor of $X$). Let $\Delta(X)\in\operatorname{CH}_{1}(X\times X)$ be the diagonal cycle. Then the correspondence
$Z=\Gamma_{\alpha}\times\Delta(X)\in\operatorname{CH}_{2}(X^{2}\times X^{2})=\operatorname{CH}_{2}(X^{3}\times X)$
@@
$\Pi_{Z}\colon\operatorname{CH}_{1}(X^{3})_{\rm hom}\to\operatorname{CH}_{0}(X)_{\rm hom};\quad\Omega\mapsto(\operatorname{pr}_{4})_{*}(\operatorname{pr}_{123}^{*}(\Omega)\cdot(\Gamma_{\alpha}\times\Delta(X))).$
@@
(4.2) $\operatorname{CH}_{1}(X^{3})_{\rm hom}$$\operatorname{CH}_{0}(X)_{\rm hom}$$J_{3}(X^{3})$$J_{1}(X),$$\scriptstyle{\Pi_{Z}}$
@@
$\Pi_{Z}(\Delta_{{\rm GKS},e}(X))=[P]+[Q]-2[\alpha(e)].$
Since for a cusp $e$, $\alpha(e)$ is also a cusp, $l\cdot\Pi_{Z}(\Delta_{{\rm GKS},e}(X))$ is non-trivial for $l\leq(p-1)/2$ by Proposition 4.1. The theorem now follows from the commutative diagram (4.2) since the Abel-Jacobi map on the right side of the diagram is an isomorphism.
@@
The automorphism group $\operatorname{Aut}(X_{3,\lambda})$ acts naturally on $\wedge^{3}H^{1}_{\rm dR}(X_{3,\lambda})$. For a subgroup $G\subset\operatorname{Aut}(X_{3,\lambda})$, we write the fixed part of $G$ for this action as $(\wedge^{3}H^{1}_{\rm dR}(X))^{G}$. Since $G_{3}=\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}\subset{\operatorname{Aut}(X_{3,\lambda})}$, we have
@@
hence by the vanishing criterion of Theorem 2.1, to prove Theorem 1.2 it suffices to show that $(\wedge^{3}H^{1}_{\rm dR}(X_{3,\lambda}))^{G_{3}}=0$. Recalling the decomposition of $H^{1}_{\rm dR}(X_{3,\lambda})$ according to the characters of $G_{3}$ from §2.3, we have
@@ 4.3. Some further remarks
Here we make a few supplementary comments on Theorem 1.1. Let $p$ be a prime number >3>3 and $e$ a cusp of $X_{p,\lambda}$. If the image of the point [P]+[Q]2[e][P]+[Q]-2[e] in one quotient curve $C_{p,\lambda}^{a,b}$ is non-torsion, then the argument in §4.1 together with Proposition 3.3 implies that $\Delta_{{\rm GKS},e^{\prime}}(X_{p,\lambda})$ is non-torsion for every choice of a based point $e^{\prime}$. We do not know how to prove that $(\varphi_{p}^{a,b})_{\ast}[P]+[Q]-2[e]$ is non-torsion for any $a,b$. However, in the cases (a,b)=(1,1),(1,p1)(a,b)=(1,1),(1,p-1) (which are the hyperelliptic cases, see Proposition 3.2), we have the following:
@@
Let $p$ be a prime number and (a,b)=(1,1),(1,p1)(a,b)=(1,1),(1,p-1). Then for any cusp $e\in X_{p,\lambda}$, the point
$(\varphi^{a,b}_{p})_{*}([P]+[Q]-2[e])\in\operatorname{Jac}(C_{p,\lambda}^{a,b})$
is torsion for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$.
@@
By Proposition 4.2, it suffices to show the case $e=c_{1,0}$. Since
@@
Acknowledgment
This article was written while the second author was visiting the University of Winnipeg. He would like to thank the Department of Mathematics and Statistics for their hospitality. The second author was supported by Waseda University Grant for Early Career Researchers (Project number: 2025E-041). This work was partially supported by the NSERC Discovery Grant “Arithmetic and Geometry of Mixed Motives”.

### v2

@@ 3.2. Properties of the cusps
Our goal in this subsection is to establish some results about the cusp points. These results will be needed in the proofs of Theorems 1.1 and 1.2, which will be given in §4.
We start with an analogue of the Manin-Drinfeld theorem for hypergeometric curves:
@@
Let $A$ and $B$ be cusps on $X_{N,\lambda}$. Then the point $A]-[B]\in\operatorname{Jac}(X_{N,\lambda})$ is a torsion point of order dividing $N^{2}$. If $\{A,B\}$ is any of $\{b_{i},c_{2,j}\}$, $\{a_{i},c_{1,j}\}$, $\{a_{i},a_{j}\}$, $\{b_{i},b_{j}\}$, $\{c_{1,i},c_{1,j}\}$ or $\{c_{2,i},c_{2,j}\}$ for any $i,j$ then [A][B][A]-[B] is annihilated by $N$.
@@
Let $K(X_{N,\lambda})$ be the canonical divisor of $X_{N,\lambda}$. For every divisor $e$ of degree $1$ on $X_{N,\lambda}$ supported on the cusps, we have
$(2(N-1)^{2}-2)e=K(X_{N,\lambda})$
@@
Thus by Proposition 3.3, $K(X_{N,\lambda})-(2(N-1)^{2}-2)e$ is torsion in the Jacobian, which proves the assertion. ∎
Our final result in this subsection is a result about the space of meromorphic functions with bounded poles at the cusps. The result may seem rather unmotivated by itself, but it will play a key role in the proof of Theorem 1.1. For a positive integer $d$ and distinct points $e_{0},\ldots,e_{n}\in X_{N,\lambda}$, let
@@
be the Riemann-Roch space of the divisor $d\displaystyle\sum_{i=0}^{n}[e_{i}]$, i.e., the vector space (over the constants $\mathbb{C}$) of meromorphic functions on $X_{N,\lambda}$ that are regular outside $\{e_{i}\mid i=0,\ldots,n\}$ and do not have a pole of order greater than $d$ at any $e_{i}$.
@@
Let $p\neq 2$ be a prime number and $l$ a positive integer such that $l\leq(p-1)/2$. Let $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and $e$ be a divisor of degree 1 on $X_{p,\lambda}$ supported on one of the four sets of cusps $\{c_{1,i}:0\leq i<p\}$, $\{c_{2,i}:0\leq i<p\}$, $\{a_{i}:0\leq i<p\}$, or $\{b_{i}:0\leq i<p\}$. Then the point
$l\cdot(\varphi^{a,b}_{p})_{*}([P]+[Q]-2e)\in\operatorname{Jac}(C_{p,\lambda}^{a,b})$
@@
We only prove the case when $e$ is supported on the $c_{1,i}$ since the other cases are proved similarly. It suffices to show that the point
$l\cdot(\varphi_{p}^{a,b})^{*}\circ(\varphi^{a,b}_{p})_{*}([P]+[Q]-2e)=l\cdot\sum_{j=0}^{p-1}\left(g_{p}^{bj,-aj}\cdot([P]+[Q])-2[c_{1,j}]\right)$
of $\operatorname{Jac}(X_{p,\lambda})$ is nontrivial. Here, the action of $g_{p}^{bj,-aj}$ in the right hand side is by pushforward of divisors. The equation uses the fact that $e$ has degree 1, as well as the fact that for every $0\leq i<p$, we have $(\varphi_{p}^{a,b})^{*}\circ(\varphi^{a,b}_{p})_{*}([c_{1,i}])=\sum_{j=0}^{p-1}[c_{1,j}]$.
Let $K$ be the function field of $X_{p,\lambda}$. Suppose that the above point of $\operatorname{Jac}(X_{p,\lambda})$ is trivial. Then there exists a function $f\in K$ such that
$\operatorname{div}(f)=l\cdot\sum_{j=0}^{p-1}\left(g_{p}^{bj,-aj}\cdot([P]+[Q])-2[c_{1,j}]\right).$
@@
It follows that $g$ has poles at $g_{p}^{0,j}(P)$ and $g_{p}^{0,j}(Q)$ for $0<j<p$ (note that these are not canceled by the zeros of $g$; also note that these are distinct from the points $c_{1,j}$ for $0\leq j<p$). But this is absurd, since $g$ is a polynomial in only $x$ and therefore it does not have any poles outside the $c_{1,j}$. ∎
We will now deduce Theorem 1.1. In fact, we may allow the base $e$ to be slightly more general:
Theorem 4.2.
Let $p\geq 3$ be a prime, $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$, and $e$ a divisor of $X_{p,\lambda}$ of degree 1 with support contained in one of the four sets of cusps $\{c_{1,i}:0\leq i<p\}$, $\{c_{2,i}:0\leq i<p\}$, $\{a_{i}:0\leq i<p\}$, or $\{b_{i}:0\leq i<p\}$. Then for every positive integer $l\leq(p-1)/2$, we have
$l\cdot\Phi_{1}(\Delta_{{\rm GKS},e}(X_{p,\lambda}))\neq 0,$
where $\Phi_{1}(\Delta_{{\rm GKS},e}(X_{p,\lambda}))$ is the complex Abel-Jacobi image of $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ in the intermediate Jacobian of the integral Hodge structure $H_{3}(X_{p,\lambda}^{3})$.
Proof.
For brevity, we set $X=X_{p,\lambda}$. We take $\Gamma_{\alpha}\in\operatorname{CH}_{1}(X\times X)$ to be the graph of $\alpha$. Let $\operatorname{pr}_{123}\colon X^{3}\times X\to X^{3}$ (resp. $\operatorname{pr}_{4}\colon X^{3}\times X\to X$) be the natural projection to the first three factors of $X$ (resp. the fourth factor of $X$). Let $\Delta\in\operatorname{CH}_{1}(X\times X)$ be the diagonal cycle. Then the correspondence
$Z=\Gamma_{\alpha}\times\Delta\in\operatorname{CH}_{2}(X^{2}\times X^{2})=\operatorname{CH}_{2}(X^{3}\times X)$
@@
$\Pi_{Z}\colon\operatorname{CH}_{1}(X^{3})_{\rm hom}\to\operatorname{CH}_{0}(X)_{\rm hom};\quad\Omega\mapsto(\operatorname{pr}_{4})_{*}(\operatorname{pr}_{123}^{*}(\Omega)\cdot(\Gamma_{\alpha}\times\Delta).$
@@
(4.2) $\operatorname{CH}_{1}(X^{3})_{\rm hom}$$\operatorname{CH}_{0}(X)_{\rm hom}$$J_{1}(X^{3})$$J_{0}(X),$$\scriptstyle{\Pi_{Z}}$
@@
$\Pi_{Z}(\Delta_{{\rm GKS},e}(X))=[P]+[Q]-2\alpha_{*}(e).$
For a divisor $e$ with support in one of the four sets $\{c_{1,i}\}$, $\{c_{2,i}\}$, $\{a_{i}\}$, or $\{b_{i}\}$, the support of $\alpha_{*}(e)$ is also in one of the same sets. Thus $l\cdot\Pi_{Z}(\Delta_{{\rm GKS},e}(X))$ is nontrivial for $l\leq(p-1)/2$ by Proposition 4.1. The theorem now follows from the commutative diagram (4.2) since the Abel-Jacobi map on the right side of the diagram is an isomorphism. ∎
@@
The automorphism group $\operatorname{Aut}(X_{3,\lambda})$ acts naturally on $\wedge^{3}H^{1}_{\rm dR}(X_{3,\lambda})$. For a subgroup $G\subset\operatorname{Aut}(X_{3,\lambda})$, we write the fixed part of $G$ for this action as $(\wedge^{3}H^{1}_{\rm dR}(X_{3,\lambda}))^{G}$. Since $G_{3}=\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}\subset{\operatorname{Aut}(X_{3,\lambda})}$, we have
@@
hence by Corollary 3.4 and the criterion of Theorem 2.1, to prove Theorem 1.2 it suffices to show that $(\wedge^{3}H^{1}_{\rm dR}(X_{3,\lambda}))^{G_{3}}=0$. Recalling the decomposition of $H^{1}_{\rm dR}(X_{3,\lambda})$ according to the characters of $G_{3}$ from §2.3, we have
@@
Remark 4.3.
For $p\geq 5$, it is not difficult to write an explicit nonzero element of $H^{3}_{\rm dR}(\operatorname{Jac}(X_{p,\lambda}))$ (in fact, an element of the primitive part) that is fixed by $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$, the automorphism $\alpha$, and the automorphism given by $(x,y)\mapsto(y,x)$.
@@ 4.3. Some further remarks
Here we make a few supplementary comments on Theorem 1.1. Let $p$ be a prime number >3>3 and $e$ a cusp of $X_{p,\lambda}$. If the image of the point [P]+[Q]2[e][P]+[Q]-2[e] in the Jacobian of one quotient curve $C_{p,\lambda}^{a,b}$ is non-torsion, then the argument in §4.1 together with Proposition 3.3 implies that $\Delta_{{\rm GKS},e^{\prime}}(X_{p,\lambda})$ is non-torsion for every choice of a degree 1 base divisor $e^{\prime}$. We do not know how to prove that $(\varphi_{p}^{a,b})_{\ast}([P]+[Q]-2[e])$ is non-torsion for any $a,b$. However, in the cases when $(a,b)$ is $(1,1)$ or (1,p1)(1,p-1) (which are the hyperelliptic cases, see Proposition 3.2), we have the following:
@@
Let $p$ be a prime number and $(a,b)\in\{(1,1),(1,p-1)\}$. Then for any divisor $e$ of degree 1 on $X_{p,\lambda}$ supported on the cusps, the point
$(\varphi^{a,b}_{p})_{*}([P]+[Q]-2e)\in\operatorname{Jac}(C_{p,\lambda}^{a,b})$
is torsion for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$. [^5] Note this it is not clear whether this statement can be deduced from the torsion-ness of the modified diagonal cycles of hyperelliptic curves. Indeed, it is not obvious whether the point $(\varphi^{a,b}_{p})_{*}([P]+[Q]-2e)$ of $\operatorname{Jac}(C_{p,\lambda}^{a,b})$ is of the form $\Pi_{Z}(\Delta_{{\rm GKS},e}(C_{p,\lambda}^{a,b}))$ for some algebraic cycle $Z$ on $(C_{p,\lambda}^{a,b})^{4}$.
@@
By Proposition 4.4, it suffices to show the case $e=[c_{1,0}]$. Since
@@
Acknowledgments
The authors would like to express sincere gratitude to the anonymous referees for their many helpful comments. This article was written while the second author was visiting the University of Winnipeg. He would like to thank the Department of Mathematics and Statistics for their hospitality. The second author was supported by Waseda University Grant for Early Career Researchers (Project number: 2025E-041). This work was partially supported by the NSERC Discovery Grant “Arithmetic and Geometry of Mixed Motives”.